OpenStudy (anonymous):

A circle has the equation 4(x-4)^2+4y^2=36 find the center and radius

OpenStudy (callisto):

Divide both sides by 4, what do you get?

OpenStudy (callisto):

@teresa65

OpenStudy (anonymous):

I am sorry but I have a hard time understanding circles

OpenStudy (callisto):

Hmm, that's just division only...

OpenStudy (anonymous):

my brain is fried- I have been at this for three hours

OpenStudy (callisto):

Simple division... \[\frac{4(x-4)^2+4y^2}{4}=\frac{36}{4}\]Simplify it, what do you get? PS: Don't worry. I had the experience of spending more than 3 hours on 1 math question. And this one is not that hard!

OpenStudy (anonymous):

Thanks!!!!

OpenStudy (callisto):

Thanks?! It's not done yet....

OpenStudy (anonymous):

sorry

OpenStudy (callisto):

You need not say sorry....... So, where are you stuck at now?

OpenStudy (anonymous):

Finding the radius and center. Algebrs was my favourite subject in high school but I have been out for 38 years

OpenStudy (callisto):

what do you get for this: \(\frac{4(x-4)^2+4y^2}{4}=\frac{36}{4}\) ?

OpenStudy (anonymous):

that is where I am stuck. I cannot seem to make any sense of this. I am having trouble with 4(x-4)^2/4

OpenStudy (anonymous):

Do i do the multiplication first then the division

OpenStudy (anonymous):

am I stupid or what? lol

OpenStudy (callisto):

Just division. \[\frac{4(x-4)^2}{4} = (x-4)^2\]

OpenStudy (anonymous):

ok now I see

OpenStudy (callisto):

Can you simplify \(\frac{4(x-4)^2+4y^2}{4}=\frac{36}{4}\) now?

OpenStudy (anonymous):

(x-4)^2 + y^2=9

OpenStudy (callisto):

Yup!!! Now, express the equation in the form \((x-h)^2 + (y-k)^2 = r^2\)

OpenStudy (anonymous):

would it be (x-2)^2 + (y-0)^2=3

OpenStudy (anonymous):

my brainis definitely gone

OpenStudy (callisto):

Not really. ^2 after 3 and things in the first bracket doesn't change. It is (x-4)^2 + (y-0)^2=3^2

OpenStudy (callisto):

Next, for a equation \((x-h)^2 + (y-k)^2 = r^2\), (h,k) are the coordinates of centre, r is the radius Can you identify the values of h, k and r?

OpenStudy (anonymous):

h=-4, k=0, and r=3?

OpenStudy (callisto):

h is 4 ... not -4 Others are correct. Now can you get the centre and radius?

OpenStudy (anonymous):

Yes I can thanks so much for the help!

OpenStudy (callisto):

Welcome!